Reduce the following equation into quadratic equation
t+2/t-2 - t-2/t+2=3/7

Answer by Boreal(15235)

t+2/t-2 - t-2/t+2=3/7
Multiply everything by (t-1)(t+2)7 to clear the fractions
7(t+2)(t+2) -7 (t-2)(t-2) =3 (t-2) (t+2) ;; (t+2)^2= t^2 + 4t + 4 ;; (t-2)^2= t^2-4t +4
7t^2+28t+28 - 7t^2+28t -28 =3t^2 -12 ;; (t-2)(t+2)=t^2-4
Left side has 56t; everything else cancels.
Equation is 56t= 3t^2-12
3t^2-56t-12=0
That is the quadratic equation